# Category-based Galaxy Image Generation via Diffusion Models

### v1

@@ 4.3 Impact of Astro-RAB on Physical Properties
Furthermore, as shown in the right panel of Fig. 6, in the Attention Fusion Unit of the U-Net bottleneck layer changes significantly. This is because the bottleneck layer serves as a critical transition point in the U-Net’s structure, bridging the preceding down-sampling stage with the subsequent up-sampling stage. The dynamic changes in the balance between attention and convolution reflect a gradual shift in feature processing from global to local. During the down-sampling stage, multiple layers of convolution and pooling operations progressively compress the input features to extract high-level information, although some details may be lost in the process. In the early units of the bottleneck layer (layers 13 and 14), the proportion of attention gradually increases, enhancing the model’s ability to capture global information through the attention mechanism. This indicates that the model is fully integrating the global features of galaxies at this stage. In contrast, in the later units of the bottleneck layer (layers 15 and 16), the proportion of attention decreases while the convolution proportion increases, signaling a transition from global information modeling to the recovery of local details. This shift effectively prepares the model for the subsequent up-sampling phase.
@@
Additionally, in our experiments, the Attention Fusion Unit utilizes an pixel window for the window attention mechanism and a convolution kernel. This configuration effectively covers the core regions of galaxies, such as the Flux within aperture photometry (4-pixel radius), which is likely one of the factors contributing to the significant enhancement of the galaxy color synthetic distance.
5 Conclusions
@@
Some galaxy generation models have been developed and are capable of generating realistic mock galaxy images, though these methods may require specific assumptions of physics models (semi-analytical models), be unable to generate high-quality images (VAEs), or struggle to perform stable generation (GANs). The most recent method in this field, Astrddpm, made use of advanced diffusion model architecture and can stably generate realistic galaxy mock images (Smith et al., 2022). The drawback of such a method, however, is that it requires significant computation resources for model training and has limited capability to benefit downstream tasks such as fueling the model training of deep learning based galaxy classification models and class-wise galaxy image segmentation models.
@@
We also evaluated the consistency between the galaxy images generated by our model and the test set in terms of color and half-light radius distribution using the Wasserstein-1 distance. Compared to Astrddpm with class embedding, GalCatDiff shows improvements of 63.46% in color metric and 66.92% in size metric, with reductions in the standard deviation of 60% and 42.86%. These results further validate the effectiveness of the Astro-RAB. Although some bias remains in the generation of minority classes and extreme samples, the model demonstrates strong overall matching in the distribution of physical metrics.
@@ Acknowledgments
HMT gratefully acknowledges the grant support from the Shuimu Tsinghua Scholar Program of Tsinghua University, the 2025 grant of the IAU Office of Astronomy for Development (IAU OAD); HMT also acknowledges support from the Department of Physics of Xi’an Jiaotong-Liverpool University, the long-lasting support from the DoA Tsinghua TAGLAB and machine learning group and JBCA machine learning group.
@@
The dataset used in this work can be accessed on request and was derived from the Galaxy Zoo DECaLS dataset (https://doi.org/10.5281/zenodo.4196266). The core code underlying this work will become accessible on GitHub after publication. We hope this code can be useful for future developers training their generative models.

### v2

@@ IV.3 Impact of Astro-RAB on Physical Properties
Furthermore, as shown in the right panel of Fig. 9, in the Attention Fusion Unit of the U-Net bottleneck layer changes significantly. This is because the bottleneck layer serves as a critical transition point in the U-Net’s structure, bridging the preceding down-sampling stage with the subsequent up-sampling stage. The dynamic changes in the balance between attention and convolution reflect a gradual shift in feature processing from global to local. During the down-sampling stage, multiple layers of convolution and pooling operations progressively compress the input features to extract high-level information, although some details may be lost in the process. In the early units of the bottleneck layer (layers 13 and 14), the proportion of attention gradually increases, enhancing the model’s ability to capture global information through the attention mechanism. This indicates that the model is fully integrating the global features of galaxies at this stage. In contrast, in the later units of the bottleneck layer (layers 15 and 16), the proportion of attention decreases while the convolution proportion increases, signaling a transition from global information modeling to the recovery of local details. This shift effectively prepares the model for the subsequent up-sampling phase. In summary, these observations suggest that global-to-local integration is accomplished primarily in the bottleneck.
To further assess the role of Astro-RAB, we conduct ablation studies on different stages of the U-Net: down-sampling, up-sampling, the bottleneck, the early bottleneck (layers 13 and 14), and the late bottleneck (layers 15 and 16). In each ablation, for the selected stage, we fix the Astro-RAB weights ( and ) to 0.5 ( = 1) and keep them constant throughout training. Table. 5 reports the evaluation metrics for all ablations, and Fig. 10 shows the per-layer values of (, ), and their ratio for the best model under each setting.
As Table. 5 indicates, removing Astro-RAB from any stage degrades both the color metrics (Color-WD) and the size metric (HLR-WD). Among all ablations, the up-sampling ablation leads to the minimal degradation, whereas ablating the bottleneck causes pronounced deterioration; in particular, removing Astro-RAB from the late bottleneck (layers 15–16) reduces performance to a level comparable to Astrddpm with class embedding. Taken together, the evidence from the Table. 5 and Fig. 9 and 10 indicates that model performance is closely associated with a global-to-local integration pattern in the bottleneck: attention weights are higher in the early bottleneck (layers 13 and 14), followed by a gradual increase in convolutional weights in the late bottleneck (layers 15 and 16). This configuration first captures global structure and then refines local details, and corresponds to closer agreement with the test-set distributions and lower WSGD. As shown in Fig. 9, the ratio exhibits a rise-then-fall trajectory across layers 13 to 16; departures from this pattern coincide with marked performance degradation.
Comparative ablations further corroborate this mechanism. In the up-sampling ablation, WSGD increases from 0.063 to 0.075 (+19%), representing the least degradation; this is plausibly because fixing ( and ) at 0.5 provides a reasonable initialization, and the bottleneck still preserves the pattern of higher attention weight in the early bottleneck (e.g., layer 14) followed by increased convolutional weight in the late bottleneck (e.g., layer 15), albeit with smaller amplitude. Ablating the late bottleneck (layers 15 and 16) yields the most severe deterioration, with WSGD increasing from 0.063 to 0.181 (+187%); in this setting, attention is suppressed in the early bottleneck and only recovers thereafter, while the expected reinforcement of convolutional weights in the late bottleneck is absent, disrupting the intended global-to-local integration. Ablating the entire bottleneck also causes substantial degradation (from 0.063 to 0.141; +124%). In the down-sampling ablation (from 0.063 to 0.096; +52%), attention briefly increases in the early bottleneck but quickly declines, failing to establish a stable progression. By comparison, ablating the early bottleneck leads to a milder decline (from 0.063 to 0.089; +41%) and smaller standard deviations across both metrics, possibly because attention increases during the latter half of down-sampling and convolutional weights strengthen in the late bottleneck, thereby partially preserving the global-to-local integration.
@@
Additionally, in our experiments, the Attention Fusion Unit utilizes an pixel window for the window attention mechanism and a convolution kernel. This configuration effectively covers the core regions of galaxies, such as the Flux within aperture photometry (1.5-pixel radius), which is likely one of the factors contributing to the significant enhancement of the weighted synthetic galaxy distance.
| Model | ||||
| GalCatDiff | ||||
| Down Stage | ||||
| Up Stage | ||||
| Bottleneck (All) | ||||
| Bottleneck (Early) | ||||
| Bottleneck (Late) |
V Conclusions
@@
Some galaxy generation models have been developed and are capable of generating realistic mock galaxy images, though these methods may require specific assumptions of physics models (semi-analytical models), be unable to generate high-quality images (VAEs), or struggle to perform stable generation (GANs). The most recent method in this field, Astrddpm, made use of advanced diffusion model architecture and can stably generate realistic galaxy mock images (Smith et al., 2022). The drawback of such a method, however, is that it requires significant computational resources for model training and has limited capability to benefit category-related downstream tasks, such as fueling the model training of deep learning based galaxy classification models and class-wise galaxy image segmentation models.
@@
We also evaluated the consistency between the galaxy images generated by our model and the test set in terms of color and half-light radius distribution using the Wasserstein-1 distance. Compared to Astrddpm with class embedding, GalCatDiff shows improvements of 49.36% in color metric and 66.92% in size metric, with reductions in the standard deviation of 16.67% and 26.32%. These results further validate the effectiveness of the Astro-RAB. Although some bias remains in the generation of minority classes and extreme samples, the model demonstrates strong overall matching in the distribution of physical metrics.
@@ Acknowledgments
We thank the anonymous reviewers for their constructive comments and suggestions. This work was supported by the National Natural Science Foundation of China (grant No. 12503111). This work was also supported by the XJTLU Research Development Fund (grant No. RDF-24-02-057). H.M.T. gratefully acknowledges support from the Department of Physics of Xi’an Jiaotong-Liverpool University, the longlasting support from the DoA Tsinghua TAGLAB and machine learning group, and the JBCA machine learning group. X.Z.F. gratefully acknowledges helpful discussions with Prof. Dennis Zaritsky, as well as assistance from Ms. Yuning Zhu in preparing the model architecture diagram.
@@
The dataset used in this work can be accessed on request and was derived from the Galaxy Zoo DECaLS dataset (Walmsley et al., 2020a). The core code underlying this work is available in Zenodo (Xingzhong, 2026). A developmental copy is availabe in GitHub666https://github.com/stevefxz/GalCatDiff/tree/main. We hope this code can be useful for future developers training their generative models.